The equilibrium constant, Kc, for the following reaction is 7.00×10-5 at 673 K.


NH4I(s) NH3(g) + HI(g)


Calculate the equilibrium concentration of HI when 0.210 moles of NH4I(s) are introduced into a 1.00 L vessel at 673 K.


[HI] =   M


1
Expert's answer
2022-02-21T17:12:03-0500


NH"_4I_{(s)}\to\>NH_{3(g)}+HI_{(g)}"


"k_c=\frac{[NH_3][HI]}{[NH_4I]}"


"7×10^{-5}=\frac{(x)(x)}{0.210-x}"


"\implies\>x^2=1.470×10^{-5}"

"x=3.834\u00d710^{-3}"


"[HI]=3.834\u00d710^{-3}M"